A fence must be installed around a circular pool at 2 m from the pool. if the pool has a radius of 4 m, determine the total cost of this fence if the fence costs $12.50 per metre to which a 15% tax must be added.

Please help! This question all has to do with circles!(:

Answer by stanbon(75887)

A fence must be installed around a circular pool at 2 m from the pool. if the pool has a radius of 4 m, determine the total cost of this fence if the fence costs $12.50 per metre to which a 15% tax must be added.
-----
Draw the picture.
You have a circle (fence) around a circle (pool).
Radius of pool = 4 meters
Radius of the fence = 6 meters.
----
Perimeter of the fence = 2*pi*r = 12(pi) meters
-----
Cost of fence: [12pi]*12.50 + 0.15[12pi*12.50]
-----
= 1.15[12pi*12.50]
= $541.92